# Shift current response in twisted double bilayer graphenes

### v1

@@ Abstract
We calculate the shift current response in twisted double bilayer graphenes (TDBG) by applying the perturbative approach to the effective continuum Hamiltonian. We have performed a systematic study of the shift current in AB-AB and AB-BA stacked TDBG, where we have investigated the dependence of the signal on the twist angle, the vertical bias voltage and the Fermi level. The numerical analyses demonstrate that a large signal is generated from the formation of the moiré minibands. Notably, we also found that there is a systematic sign reversal of the signal in the two stacking configurations below the charge neutrality point for large bias voltages. We qualitatively explain the origin of this sign reversal by studying the shift current response in AB-stacked bilayer graphene.
@@ I Introduction
Twisted double bilayer graphene (TDBG) is a moiré system where two sheets of AB-stacked bilayer graphene (BLG) are stacked with a relative twist angle, as shown in Fig. 1. A variety of studies on TDBG has been performed, such as the observation of strongly correlated phenomena [25; 26; 27; 28; 29; 30; 31; 32; 33; 34; 35; 36; 37; 38; 39; 40], owing to its highly tuneable band structure, where a band gap can be opened through the application of a vertical bias voltage. Importantly, TDBG can be fabricated in two distinct stacking structures, known as AB-AB and AB-BA, respectively shown in Fig. 1 (a) and (b). The AB-AB stacking is composed of two bilayer sheets stacked and twisted with the same orientation, while the AB-BA stacking introduces a $180^{\circ}$ offset on the second layer. A schematic diagram of the respective variants are shown in Fig. 1 (a) and (b). It is known that the two variants have completely different valley Chern numbers while hosting very similar band structures [25; 26; 21].
@@
Nowadays, the scope of investigation of the shift current has expanded to moiré materials [72; 31; 73; 74; 38; 75; 76] where there is no need to rely on the polarisation of the monolayer to observe a finite shift current; for example, two non-polar materials with a lattice mismatch can be stacked in order to create an interface with broken inversion symmetry. The shift current response in TBG, being one of the simplest moiré system, has been studied theoretically [31; 73; 77; 78]. However, the $C_{2z}$ symmetry renders the in-plane shift current to vanish [78], thus, an extra term is required in the TBG Hamiltonian to break such symmetry. This is achieved by introducing an asymmetric potential between the A and B sites which can be experimentally realised through the alignment with hBN. There have also been studies of the shift current in other graphitic systems [31; 38; 74] and heterostrucutres [63; 72; 53], such as WSe2/black phosphorene heterostructure [53] and twisted TMD heterobilayers [72; 79].
@@
In the present work, we theoretically study the shift current response in AB-AB and AB-BA stacked TDBG. This is motivated by the fact that TDBG is composed of AB-stacked BLG, which has broken $C_{2z}$ symmetry, allowing for a finite in-plane shift current response without the need for any substrate or external field unlike TBG. Furthermore, the band structure of TDBG is highly tuneable through the application of a vertical bias voltage, giving an extra degree of freedom to investigate the behaviour of the shift current. The fact that the two variants of TDBG exhibit similar band structures with distinct topologies  [25; 26; 21] offers a platform to study how this contrast influences the shift current response. One previous work has studied the shift current response in AB-AB and AB-BA stacked TDBG for fixed values of the twist angle, Fermi level and vertical bias voltage [31], however, the relationship between the two variants was unclear. Here, we systematically investigate the twist angle, Fermi level and vertical bias voltage dependence on the TDBG shift current to study its relation to the stacking configuration. In particular, we have found a systematic sign reversal in the signals of the AB-AB and AB-BA variants when a large vertical bias voltage is applied. We further give a qualitative explanation on the origin of this sign reversal by studying the shift current in AB-stacked BLG.
@@
The present paper is organised as follows. In Section II, we introduce the effective continuum Hamiltonian for TDBG and the theoretical expression of the shift current. Then, in Section III, we present our results of the shift current response in TDBG and provide a discussion on the interpretation on the results in Section IV. Finally, we will conclude the present paper in Section V.
@@ II.1 Twisted double bilayer graphene
Figure 1: (a) We show a schematic diagram of AB-stacked BLG and the stacking configuration in AB-AB stacked TDBG on the top. The corresponding moré pattern is shown in the bottom. In (b), we shows the same diagrams, but for AB-BA stacked TDBG, where the blue layer gains a $180^{\circ}$ offset.
Twisted double bilayer graphene (TDBG) is a system where two sheets of AB-stacked BLG are stacked and twisted with a relative twist angle, as shown in Fig. 1 (a), (b). To fabricate a sample of TDBG, we first stack the two BLG without twisting. There are two ways to perform this operation: one is to simply stack the two BLG sheets with the same orientation and the other is to perform a $180^{\circ}$ rotation about the $z$-axis on one sheet and then stacking the two. Then, a relative twist angle between the two BLG sheets is introduced to form a sample of TDBG. Here, the first variant is known as AB-AB stacking, while the second is known as AB-BA stacking. Their respective structures are shown in Fig. 1 (a) and (b), and the difference in their moiré patterns can be seen. Hence, the extra $180^{\circ}$ rotation leads to a different environment at the twist interface between the two stacking configurations, resulting in different band topologies [25]. It should also be noted that the AB-AB variant has the in-plane three-fold rotational symmetry $C_{3z}$ inherited from monolayer graphene and a two-fold rotational symmetry along $C_{2x}$ along the $x$-axis. Similarly, the AB-BA variant has $C_{3z}$ and $C_{2y}$ rotational symmetries.
In cases when the moiré lattice constant is much greater than the graphene lattice constant, i.e. when the twist angle is small, we can construct an effective continuum Hamiltonian for TDBG around the $K_{\pm}$ valleys. When the lower layer ($l=1$) and the upper layer ($l=2$) are rotated by an angle $\mp\theta/2$, the reciprocal lattice vectors in the respective layers are given as ${\bf\it b}^{(l)}_{i}=R\left[(-1)^{l}\theta/2\right]{\bf\it b}_{i}$, where $R(\theta)$ is the two dimensional rotation matrix. This allows us to define the moiré reciprocal lattice vectors ${\bf\it G}^{\rm{M}}_{i}={\bf\it b}^{(1)}_{i}-{\bf\it b}^{(2)}_{i}$.
@@
The velocities are defined as $v_{i}=\sqrt{3}a\absolutevalue{\gamma_{i}}/(2\hbar)$, $\gamma_{i}$ being hopping integrals, and $k_{\pm}=\xi k_{x}\pm ik_{y}$, $\xi$ being the valley index $K_{\xi}$. Namely, $\gamma_{0}=-2.4657$ eV is the intralayer nearest neighbouring hopping, $\gamma_{1}=$ 0.4 eV is the hopping between the dimerised sites, $\gamma_{3}=$ 0.32 eV and $\gamma_{4}=$ 0.044 eV are the diagonal hopping between the ($A_{1}$, $B_{2}$) and ($B_{1}$, $B_{2}$) sites, respectively. Also, ${\bf\it k}_{l}=R\left[(-1)^{l}\theta/2\right]({\bf\it k}-{\bf\it K}^{(l)}_{\xi})$ and ${\bf\it K}^{(l)}_{\xi}$ are the ${\bf\it K}_{\xi}$ valleys in the respective layers. By inspecting the Hamiltoanians, we can see that the top left block corresponds to the $4\times 4$ Hamiltonian for AB-stacked BLG. We have the same Hamiltonian in the bottom right block for $H_{\rm{AB\text{-}AB}}$, however, for $H_{\rm{AB\text{-}BA}}$, we can see that the pairs $[H({\bf\it k}_{2}),H^{\prime}({\bf\it k}_{2})]$ and $[g({\bf\it k}_{2}),g^{\dagger}({\bf\it k}_{2})]$ are interchanged, reflecting the difference in stacking configuration. The interlayer moiré potential $U({\bf\it r})$ is given by
@@
where $u=$ 0.0797 eV, $u^{\prime}=$ 0.0975 eV [25; 80] and $\omega=\exp(2\pi i/3)$ is the cube root of unity. The final term $V$ is the diagonal matrix modelling the vertical bias voltage
@@
The numerical evaluation of Hamiltonians in Eq. (1) involves expanding $U({\bf\it r})$ in term of the Fourier components with respect to ${\bf\it G}=n_{1}{\bf\it G}^{\rm{M}}_{1}+n_{2}{\bf\it G}^{\rm{M}}_{2}$, with integers $n_{1}$ and $n_{2}$. This Fourier expansion entails the introduction of a cutoff in ${\bf\it k}$ space, and in this study, we have adopted a cutoff of $\norm{{\bf\it k}}\leq 4G^{\rm{M}}$.
@@ II.2 Shift current
When two rays of light, $E_{\alpha}(\omega_{\alpha})$ and $E_{\beta}(\omega_{\beta})$ polarised in the $\alpha,~{}\beta$ direction with frequency $\omega_{\alpha},~{}\omega_{\beta}$ respectively, are shone to a sample, we expect a current density in the $\mu$ direction $j_{\mu}$ resulting from a second order nonlinear optical (NLO) response. Here, $\mu,\alpha,\beta=x,y$ in the present 2D set-up. Adopting the Einstein summation convention over repeated indices, the second order NLO response is expressed as
@@
where $\sigma^{\mu}_{\alpha\beta}$ is the second order NLO conductivity tensor and $\omega_{\Sigma}=\omega_{\alpha}+\omega_{\beta}$. It should be noted that $\sigma^{\mu}_{\alpha\beta}$ will be finite only when the system lacks inversion symmetry [48; 49], which can be checked by inspecting the inversion symmetry of both sides of Eq. (5).
@@
A close inspection of Eq. (6) allows us to deduce a couple of key features of the shift current. First, $f_{ab}\,\delta(\hbar\omega-\varepsilon_{ba})$ is the joint density of states (JDoS) which counts the number of states that are separated by energy $\hbar\omega$, and |$\absolutevalue{v^{\alpha}_{ba}}^{2}$ is the dipole matrix element which dictates the transition rules at each $k$-point. Thus, we can conclude that $a$ and $b$ respectively labels the initial and final state of the optical transition and that the magnitude of the current is proportional to the number of states available for the transition. Second, the shift current is characterised by the shift vector $R^{\mu(\alpha)}_{ba}$. It is expressed as the difference between the intraband Berry connections $A^{\mu}_{nn}$ of the initial and final bands of the optical transition, and we can understand this as being the difference in the centre-of-mass coordinate of the electron wavepacket in the initial and the final band [81; 82; 83]. The final $\partial_{\mu}\varphi^{(\alpha)}_{ba}$ term is present to ensure that $R^{\mu(\alpha)}_{ba}$ is gauge invariant.
In the present paper, we only work with Hamiltonians which are linear in ${\bf\it k}$. This allows us to rewrite Eq. (6) as [73]
@@
where we have set $q=-e$, where $e$ is the elemental charge. Here, we can view the states $a$ and $b$ as the initial and final states of the real optical transition, while state $c$ corresponds to some intermediate state that is reached through virtual transitions. The numerical evaluation of the shift current response in the present work was performed using this expression. We replace the integral over the moiré Brillouin zone by a summation over a mesh of $50\times 50$ and employ a Lorentzian function $\eta/[(\hbar\omega-\varepsilon_{ba})^{2}+\eta^{2}]$ with a broadening $\eta=$ 1 meV in place of $\delta(\hbar\omega-\varepsilon_{ba})$.
@@
Therefore, $\sigma^{\mu}_{\alpha\beta}$ only contains two independent components, $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$. Since any signal can be expressed in terms of $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$, in the subsequent sections, we solely focus on these two components. We further note that in the case where the vertical bias voltage is absent, AB-AB (AB-BA) configuration recovers the $C_{2x}$ ($C_{2y}$) rotational symmetry [25], further rendering $\sigma^{y}_{yy}$ = 0 ($\sigma^{x}_{xx}$ = 0).
@@ III Results
III.1 Intrinsic shift current response
We compute the shift current response in twisted double bilayer graphene (TDBG) at a twist angle of $\theta=0.8^{\circ}$. We specifically focus on the intrinsic signal in the absence of any bias voltage $\Delta=$ 0.

### v2

@@ Abstract
We calculate the shift current response in twisted double bilayer graphenes (TDBG) by applying the perturbative approach to the effective continuum Hamiltonian. We have performed a systematic study of the shift current in AB-AB and AB-BA stacked TDBG, where we have investigated the dependence of the signal on the twist angle, the vertical bias voltage and the Fermi level. The numerical analyses demonstrate that the signal is greatly enhanced as the twist angle is reduced. Notably, we also found that there is a systematic sign reversal of the signal in the two stacking configurations below the charge neutrality point for large bias voltages. We qualitatively explain the origin of this sign reversal by studying the shift current response in AB-stacked bilayer graphene.
@@ I Introduction
Twisted double bilayer graphene (TDBG) is a moiré system where two sheets of AB-stacked bilayer graphene (BLG) are stacked with a relative twist angle, as shown in Fig. 1. A variety of studies on TDBG has been performed, such as the observation of strongly correlated phenomena [25; 26; 27; 28; 29; 30; 31; 32; 33; 34; 35; 36; 37; 38; 39; 40], owing to its highly tuneable band structure, where a band gap can be opened through the application of a vertical bias voltage. Importantly, TDBG can be fabricated in two distinct stacking structures, known as AB-AB and AB-BA, respectively shown in Fig. 1 (a) and (b). The AB-AB stacking is composed of two bilayer sheets stacked and twisted with the same orientation, while the AB-BA stacking introduces a $180^{\circ}$ offset on the second bilayer. A schematic diagram of the respective variants are shown in Fig. 1 (a) and (b). It is known that the two variants have completely different valley Chern numbers while hosting very similar band structures [25; 26; 21].
@@
Nowadays, the scope of investigation of the shift current has expanded to moiré materials [72; 31; 73; 74; 38; 75; 76] where there is no need to rely on the polarisation of the monolayer to observe a finite shift current; for example, two non-polar materials with a lattice mismatch can be stacked in order to create an interface with broken inversion symmetry. The shift current response in TBG, being one of the simplest moiré system, has been studied theoretically [31; 73; 77; 78]. However, they generally host the $C_{2z}$ rotational symmetry, with the exception for certain commensurate stacking configurations (e.g. $D_{3}$ structure [79]), which renders the in-plane shift current to vanish [78]. To obtain a finite response, such symmetry must be broken. Approaches include introducing an asymmetric potential between the A and B sites through alignment with hBN [31; 73; 77] or taking advantage of the spontaneous symmetry breaking at low temperatures. There have also been studies of the shift current in other graphitic systems [31; 38; 74] and heterostrucutres [63; 72; 53], such as WSe2/black phosphorene heterostructure [53] and twisted TMD heterobilayers [72; 80].
@@
In the present work, we theoretically study the shift current response in AB-AB and AB-BA stacked TDBG. This is motivated by the fact that TDBG is composed of AB-stacked BLG, which has broken $C_{2z}$ symmetry, allowing for a finite in-plane shift current response without the need for any symmetry-breaking unlike TBG. Furthermore, the band structure of TDBG is highly tuneable through the application of a vertical bias voltage, giving an extra degree of freedom to investigate the behaviour of the shift current. The fact that the two variants of TDBG exhibit similar band structures with distinct topologies  [25; 26; 21] offers a platform to study how this contrast influences the shift current response. A previous study investigated the shift current response in AB-AB and AB-BA stacked TDBG at a fixed twist angle [31]. In this work, we present a systematic analysis of the shift current in TDBG over a range of twist angles, Fermi levels, and vertical bias voltages — parameter dependencies that were not addressed in the earlier study. We find that the twist angle plays an important role of enhancing the signal strength as the twist angle is reduced, owing from the decrease in the typical size of the energy gaps. We also place particular emphasis on the role of stacking configuration. Notably, we discover a systematic sign reversal in the shift current between the AB-AB and AB-BA variants under large vertical bias. We further provide a qualitative explanation for this sign reversal by analysing the shift current response in AB-stacked bilayer graphene.
@@
The present paper is organised as follows. In Section II, we introduce the effective continuum Hamiltonian for TDBG and the theoretical expression of the shift current. Then, in Section III, we present our results of the shift current response in TDBG, placing particular emphasis on the effect of the twist angle and vertical bias, and provide a discussion on the interpretation on the results in Section IV. Finally, we will conclude the present paper in Section V.
@@ II.1 Twisted double bilayer graphene
Figure 1: (a) We show a schematic diagram of AB-stacked BLG and the stacking configuration in AB-AB stacked TDBG on the top. The corresponding moiré pattern is shown in the bottom. In (b), we shows the same diagrams, but for AB-BA stacked TDBG, where the blue bilayer gains a $180^{\circ}$ offset.
Twisted double bilayer graphene (TDBG) is a system where two sheets of AB-stacked BLG are stacked and twisted with a relative twist angle, as shown in Fig. 1 (a), (b). To fabricate a sample of TDBG, we first stack the two BLG without twisting. There are two ways to perform this operation: one is to simply stack the two BLG sheets with the same orientation and the other is to perform a $180^{\circ}$ rotation about the $z$-axis on one sheet and then stacking the two. Then, a relative twist angle between the two BLG sheets is introduced to form a sample of TDBG. Here, the first variant is known as AB-AB stacking, while the second is known as AB-BA stacking. Their respective structures are shown in Fig. 1 (a) and (b), and the difference in their moiré patterns can be seen. Hence, the extra $180^{\circ}$ rotation leads to a different environment at the twist interface between the two stacking configurations, resulting in different band topologies [25]. It should also be noted that the AB-AB variant has the in-plane three-fold rotational symmetry $C_{3z}$ inherited from monolayer graphene and an out-of-plane two-fold rotational symmetry along $C_{2x}$ along the $x$-axis. Similarly, the AB-BA variant has $C_{3z}$ and $C_{2y}$ rotational symmetries.
In cases when the moiré lattice constant is much greater than the graphene lattice constant, i.e. when the twist angle is small, we can construct an effective continuum Hamiltonian for TDBG around the $K_{\pm}$ valleys. When the lower bilayer ($l=1$) and the upper bilayer ($l=2$) are rotated by an angle $\mp\theta/2$, the reciprocal lattice vectors in the respective bilayers are given as ${\bf\it b}^{(l)}_{i}=R\left[(-1)^{l}\theta/2\right]{\bf\it b}_{i}$, where $R(\theta)$ is the two dimensional rotation matrix. This allows us to define the moiré reciprocal lattice vectors ${\bf\it G}^{\rm{M}}_{i}={\bf\it b}^{(1)}_{i}-{\bf\it b}^{(2)}_{i}$.
@@
The velocities are defined as $v_{i}=\sqrt{3}a\absolutevalue{\gamma_{i}}/(2\hbar)$, $\gamma_{i}$ being hopping integrals, and $k_{\pm}=\xi k_{x}\pm ik_{y}$, $\xi$ being the valley index $K_{\xi}$. Namely, $\gamma_{0}=-2.4657$ eV is the intralayer nearest neighbouring hopping, $\gamma_{1}=$ 0.4 eV is the hopping between the dimerised sites, $\gamma_{3}=$ 0.32 eV and $\gamma_{4}=$ 0.044 eV are the diagonal hopping between the ($A_{1}$, $B_{2}$) and ($B_{1}$, $B_{2}$) sites, respectively. Also, ${\bf\it k}_{l}=R\left[(-1)^{l}\theta/2\right]({\bf\it k}-{\bf\it K}^{(l)}_{\xi})$ and ${\bf\it K}^{(l)}_{\xi}$ are the ${\bf\it K}_{\xi}$ valleys in the respective bilayers. By inspecting the Hamiltoanians, we can see that the top left block corresponds to the $4\times 4$ Hamiltonian for AB-stacked BLG. We have the same Hamiltonian in the bottom right block for $H_{\rm{AB\text{-}AB}}$, however, for $H_{\rm{AB\text{-}BA}}$, we can see that the pairs $[H({\bf\it k}_{2}),H^{\prime}({\bf\it k}_{2})]$ and $[g({\bf\it k}_{2}),g^{\dagger}({\bf\it k}_{2})]$ are interchanged, reflecting the difference in stacking configuration. The interlayer moiré potential $U({\bf\it r})$ is given by
@@
where $u=$ 0.0797 eV, $u^{\prime}=$ 0.0975 eV [25; 81] and $\omega=\exp(2\pi i/3)$ is the cube root of unity. The final term $V$ is the diagonal matrix modelling the vertical bias voltage
@@
The numerical evaluation of Hamiltonians in Eq. (1) involves expanding $U({\bf\it r})$ in term of the Fourier components with respect to ${\bf\it G}=n_{1}{\bf\it G}^{\rm{M}}_{1}+n_{2}{\bf\it G}^{\rm{M}}_{2}$, with integers $n_{1}$ and $n_{2}$. This Fourier expansion entails the introduction of a cutoff in ${\bf\it k}$ space, and we have extended the cutoff until convergence was achieved, corresponding to $\norm{{\bf\it k}}\leq 5G^{\rm{M}}$.
@@ II.2 Shift current
When two rays of light, $E_{\alpha}(\omega_{\alpha})$ and $E_{\beta}(\omega_{\beta})$ polarised in the $\alpha,~{}\beta$ direction with frequency $\omega_{\alpha},~{}\omega_{\beta}$ respectively, are shone to a sample, we expect a current density in the $\mu$ direction $j_{\mu}$ resulting from a second order nonlinear optical response (NLOR). Here, $\mu,\alpha,\beta=x,y$ in the present 2D set-up. Adopting the Einstein summation convention over repeated indices, the second order NLOR is expressed as
@@
where $\sigma^{\mu}_{\alpha\beta}$ is the second order NLOR conductivity tensor and $\omega_{\Sigma}=\omega_{\alpha}+\omega_{\beta}$. It should be noted that $\sigma^{\mu}_{\alpha\beta}$ will be finite only when the system lacks inversion symmetry [48; 49], which can be checked by inspecting the inversion symmetry of both sides of Eq. (5).
@@
A close inspection of Eq. (6) allows us to deduce a couple of key features of the shift current. First, $f_{ab}\,\delta(\hbar\omega-\varepsilon_{ba})$ is the joint density of states (JDoS) which counts the number of states that are separated in energy by $\hbar\omega$, and |$\absolutevalue{v^{\alpha}_{ba}}^{2}$ is the dipole matrix element which dictates the transition rules at each $k$-point. Thus, we can conclude that $a$ and $b$ respectively labels the initial and final state of the optical transition and that the magnitude of the current is proportional to the number of states available for the transition. Second, the shift current is characterised by the shift vector $R^{\mu(\alpha)}_{ba}$. It is expressed as the difference between the intraband Berry connections $A^{\mu}_{nn}$ of the initial and final bands of the optical transition, and we can understand this as being the difference in the centre-of-mass coordinate of the electron wavepacket in the initial and the final band [82; 83; 84]. The final $\partial_{\mu}\varphi^{(\alpha)}_{ba}$ term is present to ensure that $R^{\mu(\alpha)}_{ba}$ is gauge invariant.
In the present paper, we adopt the velocity gauge for numerical evaluations. In this gauge, the electromagnetic field is introduced via the minimal coupling approach [50] and is the preferred method when the velocity operators are known; the velocity operators can be easily calculated from Eq. (1). Furthermore, we only work with Hamiltonians which are linear in ${\bf\it k}$, which allows us to rewrite Eq. (6) as [73]
@@
where we have set $q=-e$, where $e$ is the elemental charge. Here, we can view the states $a$ and $b$ as the initial and final states of the real optical transition, while state $c$ corresponds to some intermediate state that is reached through virtual transitions. The numerical evaluation of the shift current response in the present work was performed using this expression. We replace the integral over the moiré Brillouin zone by a summation over a mesh of $50\times 50$ and employ a Lorentzian function $\eta/[(\hbar\omega-\varepsilon_{ba})^{2}+\eta^{2}]$ with a broadening $\eta=$ 1 meV in place of $\delta(\hbar\omega-\varepsilon_{ba})$. We note that this replacement may lead to a finite signal as $\omega\to 0$, where the response should strictly vanish, when the gap size $\varepsilon_{ba}$ becomes comparable to the broadening parameter $\eta$. This unphysical behaviour is an artifact of the calculation and does not affect the results for $\omega\gtrsim\eta$, where the approximation remains valid.
@@
Therefore, $\sigma^{\mu}_{\alpha\beta}$ only contains two independent components, $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$. Since any signal can be expressed in terms of $\sigma^{x}_{xx}$ and $\sigma^{y}_{yy}$, in the subsequent sections, we solely focus on these two components. We further note that in the case where the vertical bias voltage is absent, AB-AB (AB-BA) configuration recovers the out-of-plane $C_{2x}$ ($C_{2y}$) rotational symmetry [25], further rendering $\sigma^{y}_{yy}$ = 0 ($\sigma^{x}_{xx}$ = 0).
@@ III Results
In this section, we present the numerical results on the shift current response in AB-AB and AB-BA twisted double bilayer graphene (TDBG) at various twist angles $\theta$ and vertical bias voltages $\Delta$. We begin by studying how the shift current evolves as we vary $\theta$, specifically focussing on the intrinsic signal at $\Delta=$ 0. We then investigate the effect of $\Delta$ on the response, fixing $\theta=0.8^{\circ}$, and compare the resulting signals in the two variants of TDBG.
III.1 Evolution with twist angle